Any help would be appreciated!

121. Consider the circuit below. Use the node-voltage method to: a. Find the voltages at the points A,B, and C. +D b. Find the
Answer #1

su pus su ② from Circuit VA23V, Ve=su Apply kel at node B -3.+ VS 22. 12 6 Ve-3 + 2V2-1022 3V2 = 24 +13 24 VB 2 2* a 12.333V.© Piza = (Iza) = R12 (0.172) = x12 EV 2 7.24 SW: Pizn - 2 x3 =120 Pon = $161) 7 1 = (10.222) x6 (Pon 28.96 w